# LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data

### v1

@@ Abstract.
The semantic capabilities of language models (LMs) have the potential to enable rich analytics and reasoning over vast knowledge corpora. Unfortunately, existing systems lack high-level abstractions to perform semantic queries at scale. We introduce semantic operators, a declarative programming interface that extends the relational model with composable AI-based operations for semantic queries over datasets (e.g., sorting or aggregating records using natural language criteria). Each operator can be implemented and optimized in multiple ways, opening a rich space for execution plans similar to relational operators. We implement our operators and several optimizations for them in LOTUS, an open source query engine with a Pandas-like API.
We demonstrate LOTUS’ effectiveness across a series of real applications, including fact-checking, extreme multi-label classification, and search. We find that LOTUS’ programming model is highly expressive, capturing state-of-the-art query pipelines with low development overhead across these diverse applications. Specifically, on the FEVER dataset for fact-checking application, LOTUS’ programs can reproduce FacTool, recent state-of-the-art pipeline, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time. In the extreme multi-label classification task on the BioDEX dataset, LOTUS reproduces state-of-the art result quality with its join operator, while providing an efficient algorithm that runs faster than a naive join. In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods used by prior works. LOTUS is publicly available at https://github.com/stanford-futuredata/lotus.
@@ 1. Introduction
The powerful semantic capabilities of modern language models (LMs) create exciting opportunities for building AI systems that reason over vast knowledge corpora. Many applications require complex reasoning over large amounts of data, including both unstructured and structured data. For example a researcher reviewing recent ArXiv (arx, [n. d.]) preprints may want to quickly obtain a summary of relevant papers from the past week, or find the papers that report the best performance for a particular task and dataset. Similarly, a medical professional may automatically extract biomedical characteristics and candidate diagnoses from many patient reports (D’Oosterlinck et al., 2023). Likewise organizations may wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business needs and productivity (dis, [n. d.]).
@@
Each of these tasks require a form of bulk semantic processing, where the AI system must process large amounts of data in often complex query patterns to perform the reasoning task at hand. Supporting the full generality of these applications with efficient and easily programmable query systems would have transformative impact. This prospect, however, raises two important and challenging questions: first, how should developers express semantic queries, and secondly, how should we design the underlying query system to achieve high efficiency and accuracy.
Unfortunately, existing systems are insufficient for serving applications that require bulk semantic processing.
Many existing LM programming frameworks (lan, [n. d.]; lla, [n. d.])
and research works (Khattab et al., 2023, 2021; Khattab and Zaharia, 2020; Lee et al., 2019; Anantha et al., 2021; Gao et al., 2023; Izacard et al., 2022; Zelikman et al., 2022; Zhang et al., 2022) provide methods, abstractions and optimizations for retrieval-augmented generation (RAG), which first performs a semantic search over the text corpus, then invokes the LM conditioned on the user question and retrieved documents.
While RAG is one common and useful query pattern, it is limited to point lookups into the data corpus and assumes the user query can be answered by one or a small set of retrieved documents. However, bulk semantic processing pipelines may involve more complex patterns not captured by RAG, such as semantic aggregations or transformations over many documents.
Alternatively, several systems (ver, [n. d.]; dat, [n. d.]; noa, 2023a; Liu et al., 2024b) offer and optimize LLM user-defined functions (UDFs) in SQL. This model offers a logically row-wise LM-execution model for batch processing data with arbitrary prompts in composition with arbitrary SQL queries. While useful for batch-processing applications, this model cannot support LM-reasoning patterns across rows and provides a low-level of abstraction with a simple LM() function.
@@
Towards a declarative programming interface for bulk semantic processing, we propose semantic operators, which extend the relational model with AI-based operations that users can compose into powerful, reasoning-based query pipelines over structured and unstructured data. These operators include semantic filters, joins, rankings, aggregations, and projections, which take natural language expressions given by the programmer. To provide an implementation of these operators, we present the LOTUS (LLMs Over Tables of Unstructured and Structured data) system and programming model. LOTUS’ query engine efficiently executes queries with semantic operators using a variety of algorithms and optimizations for each operator, while abstracting away low-level details like model context length limits and choice of algorithms.
@@
| Performs a similarity join where each row of the source table is joined with the most semantically similar tuples from the table , using fields and as the left and right join keys, respectively. | |
@@
| Performs a projection according to the langex, returning a column with a list of substrings from the input tuples. | |
| Performs a semantic similarity clustering over column to create groups. | |
| Performs a top- search over column using query . | |
| Creates similarity index for column and saves to path. | |
@@
Figure 1 begins to demonstrate the power of LOTUS’ declarative programming model and optimized query engine. For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we can easily create 3 distinct query pipelines, each written in intuitive LOTUS programs of less than 50 lines of code by composing semantic operators (e.g. filters, maps, search and joins). The modularity and composability of these semantic operators allows us to quickly explore the design space on this task. In doing so, we find the un-optimized LOTUS programs can reproduce and improve accuracy on this task by up to compared to a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), while also providing query efficiency, by default, to maintain modest execution time.
LOTUS’ optimizer also exploits the rich implementation design space of semantic operators to leverage new algorithmic and optimization opportunities. For expensive operators like semantic joins, aggregations, ranking and filters, LOTUS implements novel algorithms that maximize parallel batched-inference opportunities, use model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) with a lightweight scoring function unique to our setting, leverage semantic similarity indices, and perform algorithmic approximations. Figure 1 demonstrates the effectiveness of these methods. Compared to the best-performing un-optimized LOTUS program, Program B shown in the figure, the logically equivalent program implemented using LOTUS’ optimizations attains higher accuracy and runs faster.
@@
We systematically evaluate LOTUS through a series of real applications, including fact-checking, extreme multi-label classification, and search. Our results show that LOTUS’ programming model is highly expressive, capturing high quality and state-of-the-art query pipelines with low development overhead for these wide-ranging applications. Specifically, on the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce a recent state-of-the-art pipeline (Chern et al., 2023), as shown in Figure 1, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time. In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that provides lower execution time than the naive algorithm, demonstrating the power of LOTUS’ declarative interface. In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods (Qin et al., 2024) used by prior works.
2. The LOTUS Programming Model
We now introduce the LOTUS programming model and show its expressive power in allowing developers to declaratively specify AI-based query pipelines that bulk process large datasets of structured and unstructured data. LOTUS extends the relational model with semantic operators, which we show in Table 1. These operators can be easily composed together with standard relational operators to build powerful programs that are transparently optimized (Section 3). In this section, we describe each semantic operator and our API for them in LOTUS, grounding each in concrete examples. While our current API implementation extends Pandas (pan, [n. d.]), LOTUS’ semantic operators could be used with other existing relational query languages and APIs, such as SQL.
@@ 2.1. Datatypes
LOTUS’ data model consists of tables with structured and unstructured text fields, and our current implementation extends Pandas (pan, [n. d.]). Figure 2 shows an example LOTUS program that loads data about ArXiv papers and performs a summarization task, which finds relevant papers, filters according to whether the paper claims to outperform a specific baseline, and then summarizes the remaining papers into a single digest. We briefly describe two core components of LOTUS programming model: its data model (Section 2.1.1), and its parameterized natural language expressions (Section 2.1.2) for specifying semantic operations over the data.
@@ 2.1.1. Data Model
LOTUS is designed to seamlessly extend the relational model.
Each row in the table, represents a logical entity. Table columns may contain either structured data or unstructured fields with free-form natural-language text. LOTUS’ semantic-relational operators can take both of these data-types as inputs. Figure 3 shows an example table, where each rows represents an ArXiv paper document, with fields for the paper’s title, ArXiv URL, abstract, ArXiv domain categories, and the publication date. These columns are then passed as parameters to semantic-relational operators, such as sem_search, sem_filter, sem_agg, as shown in Figure 2.
Additionally, LOTUS supports semantic similarity indices over natural-language text columns to provide optimized semantic query processing. These indices create embeddings over each document in the column to capture semantic similarity using embedding distance metrics. Semantic indexes can be created off-line and then loaded using sem_index and load_sem_index. Figure 2 provides an example program, where upon reading the ArXiv papers data from a CSV file, the programmer loads a semantic index over the abstract column. The program then repeatedly uses the semantically-indexed column in subsequent LOTUS operations, involving semantic search, filtering and aggregation.
@@ 2.1.2. Parameterized Natural Language Expressions (langex)
A core principle of LOTUS is to provide users with a declarative interface that separates the user-specified, logical query plan from its underlying implementation.
As such, users program with LOTUS’ semantic operators by writing parameterized natural language expressions (langex111Akin to regular expressions (regex), which specify character patterns to match in text, langex are natural language expressions to programmatically specify reasoning-based patterns over structured data and free-form text.),
rather than directly prompting an underlying LM.
Figure 2 shows an example of this, where the programmer provides a langex as the parameter to the sem_filter (line 18) and sem_agg (line 19) operations. Programmers write each langex in natural language text, parameterized by one or more data columns, which are indicated in the double brackets within the formatted string. The function of these expressions varies according to the semantic operator used and may represent a predicate, aggregation, comparator function, or projection in natural language. For instance, as shown in Figure 2, the langex signature of sem_filter provides a predicate that indicates a filter criteria to apply over paper titles and abstracts, while sem_agg takes a langex that provides an associative aggregation expression, which here indicates a many-to-one summarization task over paper abstracts.
@@
Notably, these language expressions are sufficiently versatile and easy to program with, providing an intuitive and higher-level interface to the user. This makes it simple for the programmer to specify diverse and complex, multi-step query pipelines. All operator-specific prompts are automatically handled by LOTUS’ underlying query engine, and the system can leverage existing prompt optimization techniques (Khattab et al., 2023; Yuksekgonul et al., 2024) on- or off-line.
2.2. Semantic Operators
We now overview each semantic operator and their corresponding LOTUS API. Table 1 provides a concise summary of each operator.
Sem_filter returns the subset of rows that pass the filtering condition, specified by the user’s langex. As Figure 4 shows, the langex signature provides a semantic predicate over one or more table columns and can be answered by a binary ”True” or ”False” answer.
Sem_topk ranks a set of rows according to the user-defined criteria, and returns the rows that best match the ranking criteria. The signature of the langex provides a general ranking criteria according to one or more columns. The underlying system can use this langex to impose a ranking over any subset of rows, according to the chosen implementation. As Figure 5 shows, the programmer uses the langex to specify arbitrary reasoning-based ranking criteria, such as ranking paper abstracts by the most outrageous claim made.
Programmers can also optionally specify a group-by parameter to indicate a subset of columns to group over during ranking, as shown in 6. The groupings are defined using standard equality matches over the group-by columns. To use groupings according to semantic similarity, users can also perform a sem_cluster_by and pass the resulting cluster_id column as the group-by column parameter.
Sem_join combines data from two tables, evaluating the user’s predicate to return the set of rows from the left and right table that pass. As Figure 7 shows, users specify the right join table, a langex, and optionally, the join key of the left and right tables. Here the langex contains two or more columns, and provides a predicate over the left and right tables. By default, the operator performs an inner join over the two tables, and can alternatively perform left, right or outer joins, if specified.
Sem_sim_join provides a variant of the semantic join, such that rows are matched according to their semantic similarity, rather than an arbitrary natural-language predicate. Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join. Figure 8 provides an example, where one table contains papers with an indexed column of abstracts, and the other table contains a list of research interests. The user specifies the left and right table join keys, and a parameter . The left join key may or may not be indexed, whereas the the right join key must be a semantically indexed column with its index loaded. The operator performs a left join such that for each row in the left table, the output table will contain matching rows from the right table with the highest similarity scores. The programmer can also optionally specify a return column containing the semantic similarity scores of each joined row.
Sem_agg performs an aggregation over all rows of the table. As Figure 9 shows, the langex signature provides a commutative and associative aggregation function, which can be applied over any subset of rows to produce an intermediate results. Semantic aggregations can be useful for many tasks involving many-to-one reduction patterns, such as summarization or question-answering over multiple documents. Similar to sem_topk, users can also specify a group-by parameter to use. We also provide additional flexibility, by allowing the programmer to compose semantic aggregations with sem_partition_by, which provides finer-granularity control over how documents are grouped over in each LM invocation. We describe this further below, and show its use in Section 3 for overriding the default commutativity and associativity assumptions for some tasks.
Sem_partition_by creates a row partitioning over the table, which will be used by sucessive calls to sem_agg to decide which rows to group together within each LM invocation. We find in Section 3 that this can non-trivially affect the result quality of aggregation tasks, like summarization. As Figure 10 shows, sem_partition_by takes a function, which outputs a group-id for each row. LOTUS natively supports a semantic cluster function, which takes the number of clusters to create and a column, but users can also specify arbitrary partition functions. The group-ids output by the partition function indicate to the system which rows should be grouped together, in a best-effort manner, during LM invocations. The system will aggregate over documents within each group, before merging intermediate results across groups.
Sem_map performs a natural language projection over an existing column and outputs a new column in the table. As shown in Figure 11, the user’s langex specifies what to project. This operator provides general functionality and simulates logically row-wise data-flow similar to prior works that use LLM-UDFs in relational languages (Liu et al., 2024a, c). These semantic projections are broadly useful for a variety of tasks, such as row-wise summarization, classification, and entity-extraction.
Sem_extract provides similar functionality to sem_map but provides the answers by returning a list of sub-strings from the source text. This functionality is useful for applications, such as entity extraction or fact-checking, where snippet-finding or verified quotes may be preferable to synthesized LLM-based answers. As Figure 12, the user’s langex signature specifies a projection, similar to sem_map.
Sem_index generates a semantic similarity index over the specified data column, as shown in Figure 13. To generate the semantic index, users first use the settings module to declare a retrieval model, which will be used for generating semantic embeddings and indexing the column data. The sem_index operator takes a column and a local directory path, to which the generated semantic index will be stored. As the figure shows, users can separately index multiple columns of the table.
Load_sem_index re-loads the stored semantic index upon reading the table data from disk, as shown in Figure 14. The user specifies the column corresponding to the semantic similarity index and the directory where the index was saved.
Sem_search performs a top- semantic similarity search over a semantically-indexed column, as Figure 14 shows. The user specifies the table column to search over, a query in natural language, a target of the number of results to return, and optionally indicates whether to return similarity scores as a column in the returned table.

### v2

# Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data

@@ Abstract.
The semantic capabilities of language models (LMs) have the potential to enable rich analytics and reasoning over vast knowledge corpora. Unfortunately, existing systems lack high-level abstractions to perform bulk semantic queries across large corpora. We introduce semantic operators, a declarative programming interface that extends the relational model with composable AI-based operations for bulk semantic queries (e.g., filtering, sorting, joining or aggregating records using natural language criteria). Each operator can be implemented and optimized in multiple ways, opening a rich space for execution plans similar to relational operators. We implement our operators in LOTUS, an open source query engine with a DataFrame API. Furthermore, we develop several novel optimizations that take advantage of the declarative nature of semantic operators to accelerate semantic filtering, clustering and join operators by up to while offering statistical accuracy guarantees. We demonstrate LOTUS’ effectiveness on real AI applications including fact-checking, extreme multi-label classification, and search. We show that the semantic operator model is expressive, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that achieve up to higher quality. Overall, LOTUS queries match or exceed the accuracy of state-of-the-art AI pipelines for each task while running up to 28 faster. LOTUS is publicly available at https://github.com/stanford-futuredata/lotus.
@@ 1. Introduction
The powerful semantic capabilities of modern language models (LMs) create exciting opportunities for building AI-based analytics systems that reason over vast knowledge corpora. Many applications require complex reasoning over large amounts of data, including both unstructured and structured data. For example a researcher reviewing recent ArXiv (arx, [n. d.]) preprints may want to quickly obtain a summary of relevant papers from the past week, or find the papers that report the best performance for a particular task and dataset. Similarly, a medical professional may automatically extract biomedical characteristics and candidate diagnoses from many patient reports (D’Oosterlinck et al., 2023). Likewise, organizations wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business (dis, [n. d.]).
@@
Each of these tasks require a form of bulk semantic processing, where the analytics system must process large amounts of data and apply semantic-based analysis across a whole dataset. Supporting the full generality of these applications with efficient and easy-to-use analytics systems would have a transformative impact, similar to what RDBMSes had for tabular data. This prospect, however, raises two challenging questions: first, how should developers express semantic queries, and secondly, how should we design the underlying analytics system to achieve high efficiency and accuracy.
Unfortunately, existing systems are insufficient for applications that require bulk semantic processing. Several systems provide logically row-wise LM operators, either in the form of a general-prupose LLM user-defined function (UDF) (ver, [n. d.]; dat, [n. d.]; noa, 2023; Liu et al., 2024b) or specialized row-wise operators (Liu et al., 2024c, a; Lin et al., 2024) for specific applications, such as conversational chat or data cleaning and extract-transform-load (ETL) tasks. These programming models fail to support a range of LM-based transformations across rows, such as ranking, grouping or joining records. Alternatively, many LM programming frameworks (lan, [n. d.]; lla, [n. d.]; Khattab et al., 2023) and research works (Khattab et al., 2021; Lee et al., 2019; Anantha et al., 2021; Gao et al., 2023; Izacard et al., 2022; Zelikman et al., 2022; Zhang et al., 2022) support combining LLM calls with retrieval-augmented generation (RAG) and Text2SQL, however these models are limited to point lookups to the data. RAG and Text2SQL methods first retrieve data using semantic search or LM-generated SQL calls, and then process the results with an LLM. Both are limited to analyses where an LLM can be fed a small number of documents or rows that are placed into a single model call’s context window (e.g., questions answered in one document in the corpus). However, bulk semantic processing applications may involve analysis across multiple rows, for clustering, ranking or joining an entire dataset.
@@
Towards a declarative programming interface for bulk semantic processing, we propose semantic operators, which extend the relational model with AI-based operations that users can compose into powerful, reasoning-based query pipelines over structured and unstructured data. These operators include semantic filters, joins, rankings, aggregations, and projections, which take natural language expressions given by the programmer. To provide an implementation of these operators, we present the LOTUS system, which exposes these operators through a simple DataFrame-based programming interface. LOTUS’ query engine efficiently and accurately executes queries with semantic operators, while abstracting away low-level details like model context length limits and algorithmic choices. Finally, we propose novel optimizations for the semantic filter, join, ranking and groupby operators that can improve performance by up to over a naive, expensive implementation while guaranteeing accuracy similar to the naive implementation.
@@
| Groups the tuples into categories based on the langex, which specifies a grouping criteria. | |
@@
Figure 1 begins to demonstrate the power of LOTUS’ declarative programming model and optimized query engine. For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we write an intuitive LOTUS program in less than 50 lines of code by composing semantic operators (i.e., filters, maps, and joins). The modularity and composability of these semantic operators make it easy to quickly write programs and explore the design space on this task. In doing so, we find the un-optimized LOTUS programs can reproduce and improve accuracy on this task by percentage points compared to a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), which was written in over 750 lines of code. LOTUS also provides query efficiency by default, yielding lower execution time than the FacTool implementation.
LOTUS’ optimizer exploits the rich implementation design space of semantic operators to provide novel optimizations for bulk semantic operations. To demonstrate the power of programming with semantic operators, we implement optimizations for several costly operators, i.e., semantic filter, join, ranking and group-by. Specifically, we develop a novel, lossless optimization for ranking. Additionally, for filters, joins and group-bys, we define and implement both a tractable, but expensive, reference algorithm as well as an approximation algorithms with probabilistic guarantees on accuracy with respect to the reference algorithm. Our approximation algorithms leverage build on statistical techniques used in prior works (Kang et al., 2021, 2020) with novel and efficient proxy scorers that leverage small LMs or semantic embeddings. Figure 1 demonstrates the effectiveness of these methods. The program implemented using LOTUS’ optimizations with recall and precision targets and a error probability attains high accuracy, of the unoptimized program’s accuracy, and runs faster than the unoptimized program.
We systematically evaluate the expressiveness of the semantic operator model and the efficiency of our algorithms and optimizations with statistical accuracy guarantees through several real applications, including fact-checking, extreme multi-label classification, and search, which had previously been implemented by hand in AI research papers. We show that LOTUS’ programming model is highly expressive, capturing high quality and state-of-the-art query pipelines with low development overhead for these wide-ranging applications. Furthermore, our algorithms and optimizations can speed up query execution by up to while offering statistical guarantees on accuracy.
@@
Specifically, on the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce a recent state-of-the-art pipeline (Chern et al., 2023), as shown in Figure 1, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering or lower execution time with or without batching, respectively, while providing statistical accuracy guarantees. In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that provides lower execution time than the naive algorithm, while again offering statistical guarantees on accuracy. In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods (Qin et al., 2024) used by prior works.
Our main contributions are the following:
We propose semantic operators as a programming model for AI-based analytics. Semantic operators (e.g., filters, joins, top-k ranking, aggregations, and group-bys) are declarative and can be implemented in multiple ways, creating a rich design space for execution plans.
We present efficient algorithms and query optimizations for several costly operators, i.e., semantic filter, join, top-k ranking and group-by. We develop a novel, lossless optimization for ranking and approximations for filters, joins and group-bys that provide statistical accuracy guarantees with respect to tractable, but expensive, reference algorithms.
We provide a reference implementation of semantic operators and optimizations in the open-source LOTUS system.
2. Semantic Operators
We define semantic operators as declarative transformations on one or more relations, parameterized by a natural language expression, that can be implemented by a variety of AI-based algorithms. Semantic operators provide a seamless extension to the relational model, operating over tables that may contain traditional structured data as well as unstructured fields such as free-form text. These new operators can be composed with traditional relational operators to build rich queries that can be transparently optimized, which we discuss in Section 4. Table 1 lists a core set of semantic operators we propose in this paper, which cover the common semantic transformations we observe in real-world applications and mirror the key transformation patterns in standard relational operators. Specific analytics systems may of course provide additional semantic operators or APIs that combine the ones we discuss here.
Each semantic operator takes a parameterized natural language expressions (langex for short), which are natural language expressions that specify a function over one or more attributes for the given semantic transformation. The langex signature depends on its semantic operator and may represent a predicate, aggregation, comparator function, or projection in natural language.
Figure 2 shows an example of using semantic operators in LOTUS’ dataframe API to load data about ArXiv papers and perform a summarization task, which finds relevant papers using a top-k search, filters according to whether the paper claims to outperform a specific baseline, and then summarizes the remaining papers into a single digest with an aggregation. The code in Figure 2 highlights various langex signatures: the sem_filter langex provides a predicate that indicates a filter criteria to apply over the title and abstract attributes, while sem_agg takes a langex that provides an associative aggregation expression, which here indicates a many-to-one summarization task over abstracts.
Together, each langex and semantic operator specifies a modular semantic transformation, whose behavior is defined with respect to a world model , which may be a powerful LLM that captures a probability distribution over the vocabulary .
We briefly overview several semantic operators, both ones whose definitions directly extend their analogs in traditional relational algebra, and those with more complex definitions.
Semantic Filter is a unary operator over the relation and returns the relation , where the natural language predicate is specified by the langex with respect to model . The langex is paramterized by one or more attributes from the relation that are relevant for evaluating the predicate.
Semantic Join provides a binary operator over relations and to return the relation , once again taking the langex with respect to model . Here the langex is parameterized by the left and right join keys and describes a natural language predicate over both.
Semantic Top-k imposes a ranking111This definition implies that imposes a total and consistent ordering. However, this definition can also be softened to assume partial orderings and noisy comparisons with respect to model . over the relation and returns . Here the signature of the langex provides a general ranking criteria according to one or more attributes of the input relation. The underlying system can use this langex to impose a ranking over any subset of rows, according to the chosen implementation.
Semantic Aggregation performs an aggregation over the input relation, returning . Here, the the langex signature provides a commutative and associative aggregation function, which can be applied over any subset of rows to produce an intermediate results. We note that the langex with respect to model assumes infinite context and can take an arbitrary number of input tuples. Managing finite context limits of the underlying model is an implementation detail left to the system. Semantic aggregations can be useful for many tasks involving many-to-one reduction patterns, such as summarization or question-answering over multiple documents.
Semantic Group-by takes a langex that specifies a grouping criteria and a target number of groups. As a example extending that of Figure 2, we might group-by the topics presented in each paper abstract with a target of 10 group. Here, the group-by operator must discover representative group labels and assign each tuple a label. In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n. d.]). Clustering algorithms over points in a metric space typically optimize the potential function tractably using coordinate descent algorithms, such as k-means. In the semantic group-by, the clustering task is over unstructured documents with a natural language similarity function specified by the user langex , which imposes a real-valued score between a tuple and a candidate label , and the operator poses the following optimization problem:
where are group labels, each consisting of tokens in the vocabulary, .
3. The LOTUS Programming Interface
In this section we describe the LOTUS API, which implements the semantic operator model as an extension of Pandas (pan, [n. d.]). We chose a Pandas-like API in our initial system implementation to make it easy for users to integrate LOTUS with popular AI libraries in Python. However, semantic operators could also be added to a variety of other data processing APIs and query languages, such as SQL.
@@ 3.1.1. Data Model
LOTUS’ data model consists of tables with structured and unstructured text fields. LOTUS’ semantic operators can take both of these data-types as inputs. For instance, Figure 2 shows an example LOTUS program, where the table schema over ArXiv papers consists of attributes for the paper’s title, ArXiv URL, abstract, ArXiv domain categories, and the publication date. These columns are then passed as parameters to semantic operators, such as sem_search, sem_filter, sem_agg.
Additionally, LOTUS supports semantic indices over natural-language text columns to provide optimized query processing. These indices leverage semantic embeddings over each document in the column to capture semantic similarity using embedding distance metrics. Semantic indices can be created off-line with sem_index and a specified retriever model, and then loaded using load_sem_index.
@@ 3.1.2. Parameterized Natural Language Expressions (langex)
Programmers write each langex in natural language, parameterized by one or more data columns, which are indicated in the brackets within a string, or double brackets within a formatted string, as Figure 2 shows. Notably, these language expressions are sufficiently versatile and easy to program with, providing an intuitive and higher-level interface to the user. Operator-specific LLM prompts based on these expressions are automatically constructed by LOTUS’ underlying query engine, and the underlying system can leverage existing prompt optimization techniques (Khattab et al., 2023; Yuksekgonul et al., 2024) on- or off-line.
3.2. Semantic Operators in LOTUS
We now overview the semantic operators supported in the LOTUS API, which includes the core set of operators described in Section 2, as well as several additional variants provided for convenience.
Sem_filter, Sem_join, & Sem_sim_join. The LOTUS API supports sem_filter and sem_join, both of which take a langex predicate, as described in Section 2. In addition, LOTUS provides a join variant, sem_sim_join, where tuples are matched according to their semantic similarity, rather than an arbitrary natural-language predicate. Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join, which indicates additional optimization opportunities to the query engine. Figure 3 provides an example of the sem_join compared to the sem_sim_join, where the user specifies the left and right table join keys, and a parameter .
The operator performs a left join such that for each row in the left table, the output table will contain matching rows from the right table with the highest semantically similarity scores.